# Language-Codec: Reducing the Gaps Between Discrete Codec Representation and Speech Language Models

### v3

@@ Abstract.
In recent years, large language models have achieved significant success in generative tasks (e.g., speech cloning and audio generation) related to speech, audio, music, and other signal domains. A crucial element of these models is the discrete acoustic codecs, which serve as an intermediate representation replacing the mel-spectrogram. However, there exist several gaps between discrete codecs and downstream speech language models. Specifically, 1) most codec models are trained on only 1,000 hours of data, whereas most speech language models are trained on 60,000 hours; 2) Achieving good reconstruction performance requires the utilization of numerous codebooks, which increases the burden on downstream speech language models; 3) The initial channel of the codebooks contains excessive information, making it challenging to directly generate acoustic tokens from weakly supervised signals such as text in downstream tasks. Consequently, leveraging the characteristics of speech language models, we propose Language-Codec. In the Language-Codec, we introduce a Masked Channel Residual Vector Quantization (MCRVQ) mechanism along with improved fourier transform structures, larger training datasets, refined discriminator design and optimized hyperparameter selection to address the aforementioned gaps. We compare our method with competing audio compression algorithms and observe significant outperformance across extensive evaluations. Furthermore, we also validate the efficiency of the Language-Codec on downstream speech language models. The source code and pretrained models can be accessed at https://github.com/jishengpeng/languagecodec.
discrete codec, speech language models, speech representation learning, residual vector quantization
ccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Modeling and simulationccs: Computing methodologies Machine learning
1. Introduction
@@
In recent times, significant achievements have been made by large-scale language models (Brown et al., 2020) in generative tasks involving such as multiple speaker speech syntheses (Wang et al., 2023; Kharitonov et al., 2023; Ji et al., 2023), music generation (Agostinelli et al., 2023), and audio generation (Kreuk et al., 2022). This success can largely be attributed to the utilization of discrete acoustic codec representations produced by neural codec models (Zeghidour et al., 2021; Défossez et al., 2022), which enable powerful transformer-based sequence-to-sequence modeling approaches for audio generation. The primary objective of discrete codec models is to convert a high-resolution audio signal (e.g., audio sampled at 44 kHz per second) into the two-dimensional discrete space. This transformation allows for the maximal compression of the speech signal in the time and frequency domains while maintaining excellent audio reconstruction quality.
@@
Currently, most end-to-end discrete codec models (Zeghidour et al., 2021; Défossez et al., 2022; Yang et al., 2023) typically adopt a three-stage structure consisting of an encoder, a Residual Vector Quantization (RVQ) module, and a decoder. The encoder performs downsampling of the audio signal in the time domain to obtain compressed audio frames. Each compressed audio frame is then quantized by a series of quantizers, with each quantizer operating on the residual of the previous one. The number of quantizers determines the overall bitrate. The decoder, on the other hand, performs upsampling in the time domain to reconstruct the audio signal from the quantizer outputs.
@@
While most codec models strive to optimize their architecture (Kumar et al., 2023; Défossez et al., 2022; Siuzdak, 2023), resulting in satisfactory audio reconstruction quality, there are still areas worth investigating in the construction of a discrete acoustic codec space that facilitates downstream speech language model modeling (Wang et al., 2023; Borsos et al., 2023; Kharitonov et al., 2023). Specifically, we believe that there exist gaps between discrete codec models and speech language models, which can be characterized as follows: 1) Currently, the majority of training data for mainstream codec models consists of approximately one thousand hours (Panayotov et al., 2015), whereas downstream speech language models have already attempted training on tens or even hundreds of thousands (Kahn et al., 2020) of speech samples. For codec models to exhibit robustness and generalizability across multiple languages and real-world scenarios, larger training datasets are still required; 2) In order to generate high-quality audio through token modeling with the neural codecs, the rate of discrete representation must be increased, which leads to either exponential growth in codebook size or the generation of long token sequences. However, there is a need for fewer codebooks to achieve this goal; 3) We have discovered that a single channel of codebook is sufficient to reconstruct a significant portion of the audio signal. The RVQ structure, in particular, results in the first channel of the codebook containing excessive information. Consequently, in downstream tasks, whether unconditionally or based on weak conditioning such as text, efficiently generating long and high-quality audio segments remains an unresolved challenge.
@@
Based on the findings mentioned above, we attempted to construct a discrete codec model that is more suitable for downstream speech language models. Our objective is to include less information in the first channel of the codebook while increasing the missing information on limited channels. Therefore, we devised the Masked Channel Residual Vector Quantization (MCRVQ) mechanism, which employs the masking mechanism to restrict the quantizers of the first three channels to learn only the compressed audio frame information in the specified space. Simultaneously, leveraging an improved Fourier transform structure (Siuzdak, 2023), the largest training dataset, and optimized hyperparameter selection, Language-Codec achieves excellent reconstruction quality on various test datasets using only four channels, thereby enhancing its compatibility with downstream models. The contributions of Language-Codec are as follows:
@@
  • By utilizing the largest speech training data, an enhanced decoder upsampling structure and hyperparameter selection. Language-Codec achieves excellent audio reconstruction quality with only four codebook channels.
@@
  • The code, pre-trained models and demo of Language-Codec have already been open-source.
2. Related Work
2.1. Discrete Codec Model
@@
In recent times, neural acoustic codecs (Zeghidour et al., 2021; Défossez et al., 2022; Yang et al., 2023; Siuzdak, 2023; Zhang et al., 2023a; Kumar et al., 2024; Du et al., 2023) have demonstrated remarkable capabilities in reconstructing high-quality audio at extremely low bitrates. Consequently, these codecs have facilitated the application of discrete modeling to a wide range of audio signals, including zero-shot TTS, music generation, and audio generation. Typically, these methods employ an encoder to extract deep features in a latent space, which are subsequently quantized before being fed into the decoder.
@@
To elaborate, Soundstream (Zeghidour et al., 2021) utilizes a model architecture comprising a fully convolutional encoder-decoder network and a residual vector quantizer (RVQ) to effectively compress speech. Encodec (Défossez et al., 2022) employs a streaming encoder-decoder architecture with a quantized latent space, trained in an end-to-end fashion. HiFi-Codec (Yang et al., 2023) introduces a group-residual vector quantization (GRVQ) technique to reduce the number of quantizers. Vocos (Siuzdak, 2023) aims to bridge the gap between time-domain and Fourier-based neural vocoders for high-quality audio synthesis. In order to narrow the gaps between text and acoustic codec tokens, SpeechTokensizer (Zhang et al., 2023a) introduces the concept of using semantic tokens in the first channel of discrete codecs. This transition helps mitigate the disparity between text and acoustic tokens. The Language-Codec adopts a similar model structure as the aforementioned methods, while additionally employing the MCRVQ mechanism to evenly distribute information from the first three channels, thereby reducing the gap between text and acoustic tokens. Furthermore, the Language-Codec leverages larger training datasets, improved model structures, enhanced loss function, and better hyperparameter selection to enhance audio reconstruction with fewer channels.
2.2. Zero-shot TTS
Zero-shot speech synthesis, also known as voice cloning, refers to the capability of synthesizing the voice of an unseen speaker using only a brief audio prompt. In recent months, significant progress has been made in zero-shot TTS, thanks to advancements in discrete acoustic codec representations and generative large-scale models. Several notable works have emerged as a result. One such work is VALL-E (Wang et al., 2023), which utilizes discrete codec representations and employs a cascaded approach by combining autoregressive and non-autoregressive models. This framework effectively preserves the powerful contextual capabilities of language models and can clone the voice of a target speaker with just a 3-second audio prompt. Building upon the cascaded structure of VALL-E, VALL-E X (Zhang et al., 2023b) extends zero-shot TTS to multiple languages. Another noteworthy approach is NaturalSpeech2 (Shen et al., 2023), which utilizes continuous vectors instead of discrete neural codec tokens. It introduces in-context learning to a latent diffusion model, enhancing the synthesis quality. Megatts(Jiang et al., 2023b, a) utilizes traditional mel-spectrograms, decoupling timbre and prosody and further modeling the prosody using an autoregressive approach. SpearTTS (Kharitonov et al., 2023) and Make-a-Voice (Huang et al., 2023) employ semantic tokens to bridge the gap between text and acoustic features. They achieve this by decoupling timbre and prosody and further modeling the prosody using an autoregressive approach. MobileSpeech (Ji et al., 2024), on the other hand, employs a masking mechanism to parallelly generate the first channel and the remaining channels of the discrete acoustic codecs in the SMD structure. In our downstream experiments, we select VALL-E (Wang et al., 2023) and MobileSpeech (Ji et al., 2024) as representatives of autoregressive and non-autoregressive models, respectively, to evaluate the performance of the discrete acoustic codecs in zero-shot TTS tasks.
@@ 2.2. Zero-shot TTS
Figure 1. The overall architecture for Language-Codec. On the far left is the encoder downsampling module, which still utilizes the model structure of Encodec. On the far right is the decoder upsampling module, where we have replaced it with Vocos’ model structure. the middle part is the Masked Channel Residual Vector Quantization module, with the gray blocks indicating the masked portion of temporal information.
3. Language-Codec
@@
In this section, we will first introduce the overall structure of Language-Codec, followed by a detailed focus on the encoder module, decoder module, and Masked Channel Residual Vector Quantization module. Furthermore, we will proceed to elaborate on the specific intricacies of Language-Codec’s training processes, with the explicit details of the loss and discriminator.
3.1. Overall
@@
(1) $X=\left\{x_{1},x_{2},\cdots,x_{T}\right\},T=d*sr$
Where $sr$ is the audio sample rate and $d$ is the audio duration. Language-Codec passes the raw audio $X$ through three modules. 1) an encoder network that takes the input audio and generates a latent feature representation $Z$; 2) a group-residual quantization layer $q$ that produces a compressed representation $Z_{q}$; and 3) a decoder that reconstructs the audio signal $\tilde{X}$ from the compressed latent representation $Z_{q}$. The model is trained end-to-end, optimizing a reconstruction loss applied over both time and frequency domains, along with a perceptual loss in the form of discriminators operating at different resolutions.
3.2. Encoder and Decoder
Follow Encodec (Défossez et al., 2022), the encoder model consists of a 1D convolution with $C$ channels and a kernel size of 7 followed by $B$ convolution blocks. Each convolution block is composed of a single residual unit followed by a down-sampling layer consisting of a stridden convolution, with a kernel size of twice the stride $S$. The residual unit contains two convolutions with kernel size 3 and a skip-connection. The number of channels is doubled whenever down-sampling occurs. The convolution blocks are followed by a two-layer LSTM for sequence modeling and a final 1D convolution layer with a kernel size of 7 and $D$ output channels. Following Encodec (Défossez et al., 2022), we use $C$ = 32, $B$ = 4, and (2, 4, 5, 8) as $S$. We use ELU as a non-linear activation function. With this setup, Language-Codec outputs 75 latent steps per second of audio at 24 kHz.
Language-Codec does not employ a mirrored decoder upsampling structure. Instead, following Vocos (Siuzdak, 2023), we maintain consistent feature resolution at all depths, achieving waveform upsampling through inverse Fourier transform. In the decoder section, the target audio signal $\tilde{X}$ is represented using Short-Time Fourier Transform (STFT):
(2) $STFT(\tilde{X}_{\left[m,k\right]})=\sum_{n=0}^{N}\tilde{X}\left[n\right]w\left[n-m\right]e^{-j2\pi kn/K}$
@@ 3.2. Encoder and Decoder
Therefore, for the representation of the intermediate signals $Z_{q}$ after quantization, the Language-Codec only needs to input $Z_{q}$ into the ConvNeXt (Liu et al., 2022), which serves as the fundamental backbone. Subsequently, a Fourier transform is performed on the real-valued signals. In ConvNeXt Block, it first embeds the input features into a hidden dimensionality and then applies a sequence of convolutional blocks. Each block is composed of a large-kernel-sized depthwise convolution, followed by an inverted bottleneck that projects features into a higher dimensionality using pointwise convolution. GELU (Gaussian Error Linear Unit) activations are used within the bottleneck, and Layer Normalization is employed between the blocks. Regarding the transformation of real-valued signals, we utilize a single side band spectrum, resulting in $n_{fft}/2+1$ coefficients per frame. Since we parameterize the model to output both phase and magnitude values, the activations of the hidden dimensions are projected into a tensor $h$ with $n_{fft}+2$ channels and subsequently split into:
(3) $q=h\left[1:n_{fft}/2+1\right];p=h\left[n_{fft}/2+2:n\right]$
@@
3.3. Masked Channel Residual Vector Quantization
@@
(7) $P(Z_{q}|Z)=\prod_{i=1}^{N_{q}}P(\hat{Z_{i}}|\bar{M}Z)P(\hat{Z}_{N_{q}+1}|Z-\sum_{i=1}^{N_{q}}\hat{Z}_{i})\prod_{j=N_{q}+1}^{N-1}P(\hat{Z}_{j+1}|Z-\hat{Z}_{j})$
3.4. Discriminator and Loss
The adversarial loss is used to promote perceptual quality. We employ the multi-period discriminator (MPD) as defined by (Kong et al., 2020) and multi-resolution discriminator (MRD) (Jang et al., 2021). Furthermore, to learn discriminative features about a specific sub-band and provide a stronger gradient signal to the generator, following (Kumar et al., 2024), we use a multi-scale discriminator (MSD) and a complex STFT discriminator (Zeghidour et al., 2021) at multiple time-scales (Défossez et al., 2022). We adopt a hinge loss formulation instead of the least squares GAN objective, as suggested by (Zeghidour et al., 2021). To train the discriminator, we can optimize the following objective function $\mathcal{L}_{dis}(X,\tilde{X})$:
(8) $\frac{1}{K}\sum_{k=1}^{K}max(0,1-D_{k}(X))+max(0,1+D_{k}(\tilde{X}))$
The variable $K$ represents the number of discriminators. $D$ represents the $k$-th discriminator. Regarding the loss for the generator, the Language-Codec model consists of four components: quantizer loss, mel-spectrum reconstruction loss, adversarial loss, and feature matching loss. The quantizer loss can be defined as follows:
(9) $\mathcal{L}_{q}(Z,Z_{q})=\sum_{i=1}^{N}\left\|Z_{i}-\hat{Z}_{i}\right\|_{2}^{2}$

### v4

# Language-Codec: Bridging Discrete Codec Representations and Speech Language Models

@@ Abstract
In recent years, large language models have achieved significant success in generative tasks related to speech, audio, music, and other signal domains. A crucial element of these models is the discrete acoustic codecs, which serve as an intermediate representation replacing the mel-spectrogram. However, there exist several gaps between discrete codecs and downstream speech language models. Specifically, 1) Due to the reconstruction paradigm of the Codec model and the structure of residual vector quantization, the initial channel of the codebooks contains excessive information, making it challenging to directly generate acoustic tokens from weakly supervised signals such as text in downstream tasks. 2) numerous codebooks increases the burden on downstream speech language models. Consequently, leveraging the characteristics of speech language models, we propose Language-Codec. In the Language-Codec, we introduce a Masked Channel Residual Vector Quantization (MCRVQ) mechanism along with improved fourier transform structures and attention blocks, refined discriminator design to address the aforementioned gaps. We compare our method with competing audio compression algorithms and observe significant outperformance across extensive evaluations. Furthermore, we also validate the efficiency of the Language-Codec on downstream speech language models. Codes are available at https://github.com/jishengpeng/Languagecodec.
Language-Codec: Bridging Discrete Codec Representations and Speech Language Models
Shengpeng Ji1thanks: Equal contribution., Minghui Fang [^1] footnotemark: 1, Jialong Zuo [^1] footnotemark: 1, Ziyue Jiang1, Dingdong Wang Hanting Wang1, Hai Huang1, Zhou Zhao1thanks: Corresponding author. 1 Zhejiang University
1 Introduction
@@
In recent times, significant achievements have been made by large-scale language models Brown et al. (2020) in generative tasks involving such as multiple speaker speech syntheses Wang et al. (2023); Kharitonov et al. (2023); Ji et al. (2023), music generation Agostinelli et al. (2023), and audio generation Kreuk et al. (2022). This success can largely be attributed to the utilization of discrete acoustic codec representations produced by neural codec models Zeghidour et al. (2021); Défossez et al. (2022), which enable powerful transformer-based sequence-to-sequence modeling approaches for audio generation. The primary objective of discrete codec models is to convert a high-resolution audio signal (e.g., audio sampled at 44 kHz per second) into the two-dimensional discrete space. This transformation allows for the maximal compression of the speech signal in the time and frequency domains while maintaining excellent audio reconstruction quality.
@@
Currently, most end-to-end discrete codec models Zeghidour et al. (2021); Défossez et al. (2022); Yang et al. (2023) typically adopt a three-stage structure consisting of an encoder, a Residual Vector Quantization (RVQ) module, and a decoder. The encoder performs downsampling of the audio signal in the time domain to obtain compressed audio frames. Each compressed audio frame is then quantized by a series of quantizers, with each quantizer operating on the residual of the previous one. The number of quantizers determines the overall bitrate. The decoder, on the other hand, performs upsampling in the time domain to reconstruct the audio signal from the quantizer outputs.
@@
While most codec models strive to optimize their architecture Kumar et al. (2023); Défossez et al. (2022); Siuzdak (2023), resulting in satisfactory audio reconstruction quality, there are still areas worth investigating in the construction of a discrete acoustic codec space that facilitates downstream speech language model modeling Wang et al. (2023); Borsos et al. (2023); Kharitonov et al. (2023). Specifically, we believe that there exist gaps between discrete codec models and speech language models, which can be characterized as follows: 1) The Codec model is inherently designed for information compression; therefore, the training objective of the Codec model aims to preserve as much information as possible within the codebook space to enhance reconstruction. We have discovered that a single channel of codebook is sufficient to reconstruct a significant portion of the audio signal. The RVQ structure, in particular, results in the first channel of the codebook containing excessive information. Consequently, in downstream tasks, whether unconditionally or based on weak conditioning such as text, efficiently generating long and high-quality audio segments remains an unresolved challenge. 2) In order to generate high-quality audio through token modeling with the neural codecs, the rate of discrete representation must be increased, which leads to either exponential growth in codebook size or the generation of long token sequences. Therefore, there is a need for fewer codebooks to achieve this goal.
@@
Based on the findings mentioned above, we attempted to construct a discrete codec model that is more suitable for downstream speech language models. Our objective is to include less information in the first channel of the codebook while increasing the missing information on limited channels. We consider that within downstream speech language models, the first-layer quantizer of the Codec model serves as an intermediary module bridging textual input and subsequent quantizers. By judiciously reducing information within the first-layer quantizer, employing text (which inherently carries less information compared to speech) to generate first Codec(codec in the first quantizer) with lower information content can be more easy. Therefore, we devised the Masked Channel Residual Vector Quantization (MCRVQ) mechanism, which employs the masking mechanism to restrict the quantizers of the first three channels to learn only the compressed audio frame information in the specified space. Simultaneously, We designed a new, more powerful decoder with improved Fourier transform structure Siuzdak (2023) and attention block. We alos add a complex STFT discriminator Kumar et al. (2024a) at multiple time-scales Défossez et al. (2022). Through these modules with enhanced sampling and reconstruction capabilities, Language-Codec achieves excellent reconstruction quality on various test datasets using only four channels, thereby enhancing its compatibility with downstream models. The contributions of Language-Codec are as follows:
@@
  • By utilizing a modern, powerful decoder and multi-scale discriminator. Language-Codec achieves excellent audio reconstruction quality with only four codebook channels.
@@
  • The code and pre-trained models of Language-Codec will be open-source.
2 Related Works
@@
In recent times, neural acoustic codecs Zeghidour et al. (2021); Défossez et al. (2022); Yang et al. (2023); Siuzdak (2023); Zhang et al. (2023); Kumar et al. (2024b); Du et al. (2023); Ji et al. (2024b) have demonstrated remarkable capabilities in reconstructing high-quality audio at extremely low bitrates. Consequently, these codecs have facilitated the application of discrete modeling to a wide range of audio signals, including zero-shot TTS, music generation, and audio generation. Typically, these methods employ an encoder to extract deep features in a latent space, which are subsequently quantized before being fed into the decoder.
@@
To elaborate, Soundstream Zeghidour et al. (2021) utilizes a model architecture comprising a fully convolutional encoder-decoder network and a residual vector quantizer (RVQ) to effectively compress speech. Encodec Défossez et al. (2022) employs a streaming encoder-decoder architecture with a quantized latent space, trained in an end-to-end fashion. HiFi-Codec Yang et al. (2023) introduces a group-residual vector quantization (GRVQ) technique to reduce the number of quantizers. Vocos Siuzdak (2023) aims to bridge the gap between time-domain and Fourier-based neural vocoders for high-quality audio synthesis. In order to narrow the gaps between text and acoustic codec tokens, HILCodec Ahn et al. (2024) introduces the MFBD discriminator to guide codec modeling. APCodec Ai et al. (2024) further enhances reconstruction quality by incorporating ConvNextV2 modules in the encoder and decoder. SpeechTokensizer Zhang et al. (2023) introduces the concept of using semantic tokens in the first channel of discrete codecs. Recently, both SemantiCodec Liu et al. (2024) and DAC Kumar et al. (2024a) have demonstrated strong reconstruction performance. SemantiCodec employs a VQ-VAE framework and utilizes a diffusion model to model the VAE’s latent space. While it achieves excellent reconstruction at low bitrates, its inference speed is approximately 100 times slower than that of standard codec models. DAC Kumar et al. (2024a) incorporates strategies such as factorized codes, L2-normalized codes, and quantizer dropout, making it one of the most advanced codec models.
The primary distinction from the previously mentioned models lies in the fact that, while Language-Codec aims to optimize codec reconstruction performance, it places greater emphasis on the gap between the reconstruction paradigm and downstream generative models. We hypothesize that the first-layer quantizer serves as an intermediary module, connecting the textual input with subsequent quantizers in speech language models. To address this challenge, we propose the MCRVQ mechanism as a solution within the Language-Codec framework.
@@
Figure 1: The overall architecture for Language-Codec. On the far left is the encoder downsampling module, which still utilizes the model structure of Encodec. On the far right is the decoder upsampling module, where we have replaced it with Vocos’ model structure. The middle part is the Masked Channel Residual Vector Quantization module, with the gray blocks indicating the masked portion of temporal information. The dashed lines within the MCRVQ module indicate that the corresponding representations exhibit a decrease in residual values.
3 Language-Codec
@@
In this section, we will first introduce the overall architecture of Language-Codec, followed by a detailed focus on the encoder module, new decoder module, and Masked Channel Residual Vector Quantization module. Furthermore, we will proceed to elaborate on the specific intricacies of Language-Codec’s training processes, with the explicit details of the loss and discriminator.
3.1 Overall
@@
$X=\left\{x_{1},x_{2},\cdots,x_{T}\right\},T=d*sr$ (1)
Where $sr$ is the audio sample rate and $d$ is the audio duration. Language-Codec passes the raw audio $X$ through three modules. 1) an encoder network that takes the input audio and generates a latent feature representation $Z$; 2) a combination of parallel and serialized quantization layer $q$ that produces a compressed representation $Z_{q}$; and 3) a decoder that reconstructs the audio signal $\tilde{X}$ from the compressed latent representation $Z_{q}$. The model is trained end-to-end, optimizing a reconstruction loss applied over both time and frequency domains, along with a perceptual loss in the form of discriminators operating at different resolutions.
3.2 Encoder and Decoder
Follow Encodec Défossez et al. (2022), the encoder model consists of a 1D convolution with $C$ channels and a kernel size of 7 followed by $B$ convolution blocks. Each convolution block is composed of a single residual unit followed by a down-sampling layer consisting of a stridden convolution, with a kernel size of twice the stride $S$. The residual unit contains two convolutions with kernel size 3 and a skip-connection. The number of channels is doubled whenever down-sampling occurs. The convolution blocks are followed by a two-layer LSTM for sequence modeling and a final 1D convolution layer with a kernel size of 7 and $D$ output channels. Following Encodec Défossez et al. (2022), we use $C$ = 32, $B$ = 4, and (2, 4, 5, 8) as $S$. We use ELU as a non-linear activation function. With this setup, Language-Codec outputs 75 latent steps per second of audio at 24 kHz.
Language-Codec does not employ a mirrored decoder upsampling structure. The standard practice involves using a stack of dilated convolutions to increase the receptive field, and transposed convolutions to sequentially upsample the feature sequence to the waveform. However, this design is known to be susceptible to aliasing artifactsInstead, following Vocos Siuzdak (2023), we maintain consistent feature resolution at all depths, achieving waveform upsampling through inverse Fourier transform. In the decoder section, the target audio signal $\tilde{X}$ is represented using Short-Time Fourier Transform (STFT):
$STFT(\tilde{X}_{\left[m,k\right]})=\sum_{n=0}^{N}\tilde{X}\left[n\right]w\left[n-m\right]e^{-j2\pi kn/K}$ (2)
@@ 3.2 Encoder and Decoder
Therefore, for the representation of the intermediate signals $Z_{q}$ after quantization, the Language-Codec only needs to input $Z_{q}$ into the conv1D layer, attention block, ConvNeXt Liu et al. (2022) blocks, which serves as the fundamental backbone. Subsequently, a Fourier transform is performed on the real-valued signals. Notably, we introduced an attention module in the decoder to enhance the sequence modeling capability of the upsampling module. During experiments, we observed that although 3-second audio segments were randomly selected during training, there were no issues with length extrapolation when reconstructing longer audio segments during inference. In ConvNeXt Block, it first embeds the input features into a hidden dimensionality and then applies a sequence of convolutional blocks. Each block is composed of a large-kernel-sized depthwise convolution, followed by an inverted bottleneck that projects features into a higher dimensionality using pointwise convolution. GELU (Gaussian Error Linear Unit) activations are used within the bottleneck, and Layer Normalization is employed between the blocks. Regarding the transformation of real-valued signals, we utilize a single side band spectrum, resulting in $n_{fft}/2+1$ coefficients per frame. Since we parameterize the model to output both phase and magnitude values, the activations of the hidden dimensions are projected into a tensor $h$ with $n_{fft}+2$ channels and subsequently split into:
$q=h\left[1:n_{fft}/2+1\right];p=h\left[n_{fft}/2+2:n\right]$ (3)
@@
3.3 Masked Channel Residual Vector Quantization
@@
$P(Z_{q}|Z)$ $=\prod_{i=1}^{N_{q}}P(\hat{Z_{i}}|\bar{M}Z)P(\hat{Z}_{N_{q}+1}|Z-\sum_{i=1}^{N_{q}}\hat{Z}_{i})$ (7)
×$\times\prod_{j=N_{q}+1}^{N-1}P(\hat{Z}_{j+1}|Z-\hat{Z}_{j})$
3.4 Discriminator and Loss
The adversarial loss is used to promote perceptual quality. We employ the multi-period discriminator (MPD) as defined by Kong et al. (2020) and multi-resolution discriminator (MRD) Jang et al. (2021). Furthermore, to learn discriminative features about a specific sub-band and provide a stronger gradient signal to the generator, following Kumar et al. (2024b), we use a multi-scale discriminator (MSD) and a complex STFT discriminator Zeghidour et al. (2021) at multiple time-scales Défossez et al. (2022). We adopt a hinge loss formulation instead of the least squares GAN objective, as suggested by Zeghidour et al. (2021). To train the discriminator, we can optimize the following objective function $\mathcal{L}_{dis}(X,\tilde{X})$:
$\frac{1}{K}\sum_{k=1}^{K}max(0,1-D_{k}(X))+max(0,1+D_{k}(\tilde{X}))$ (8)
The variable $K$ represents the number of discriminators. $D_{k}$ represents the $k$-th discriminator. Regarding the loss for the generator, the Language-Codec model consists of four components: quantizer loss, mel-spectrum reconstruction loss, adversarial loss, and feature matching loss. The quantizer loss can be defined as follows:
$\mathcal{L}_{q}(Z,Z_{q})=\sum_{i=1}^{N}\left\|Z_{i}-\hat{Z}_{i}\right\|_{2}^{2}$ (9)